Show the COMPLETE SOLUTION for the given problem.

1. Use Lagrange Multiplier to determine the dimensions of a rectangular box, open at the top, having a volume of 32 cubic feet and requiring the least amount of material for its construction.
1
Expert's answer
2020-03-26T12:29:43-0400

"\text{We label the dimensions as }l,b,h.\\ \text{Our objective function is of the form }\\\mathrm{min}\; L_0=lb+2bh+2lh \\
\text{Our constraint is on the volume }lbh=32\\
\text{Thus, our unconstrained objective function is }\\
L=(lb+2bh+2lh)-\lambda(32-lbh)\\
\frac{\partial L}{\partial l}=(b+2h)+\lambda\\
\frac{\partial L}{\partial b}=(l+2h)+\lambda\\
\frac{\partial L}{\partial h}=2(b+l)+\lambda\\
\frac{\partial L}{\partial \lambda}=32-lbh\\
\text{Setting all the partial derivatives to 0, we get }\\
2(l+b)=(b+2h)=(l+2h)\implies l=b=\frac{2}{3}h\\
\text{Since }lbh=32, \; l=b=\frac{4}{3^{1/3}}, \; h=2*3^{2/3}\\
\text{These are the optimum dimensions}"